# Teacher Encoder-Student Decoder Denoising Guided Segmentation Network for Anomaly Detection

### v3

@@ 3.2. Parallel Convolutional Attention Recalibration Module
In ResNet18, each of the four convolutional layers (i.e., conv2_x, conv3_x, conv4_x, conv5_x) consists of two residual blocks. We propose a novel Parallel Convolutional Attention Recalibration (PCAR) Module. By using parallel convolutions to extract multi-scale spatial information, the PCAR module can accurately localize and segment anomalous regions while effectively eliminating irrelevant noise present in the detection images. We specifically use the PCAR module to replace the second 3x3 convolution in the residual block of the original ResNet18 architecture. The overall architecture of the PCAR module is shown in Fig. 4.
@@
First, the PCAR module receives the output feature $F\in\mathbb{R}^{H\times W\times C}$ from the first $3\times 3$ convolution layer of the Residual Block, where $H$, $W$, and $C$ represent the height, width, and number of channels of the feature map, respectively. The input feature $F$ is then passed through three parallel $3\times 3$ convolutions, each of which independently extracts features from different perspectives, resulting in different feature maps $F_{i}$ (where $i=1,2,3$), thus enhancing the diversity of the features. The convolution operation is represented as:
(5) $F_{i}=\text{Conv}{3\times 3}(F)$
Next, the output $F_{i}$ from each convolution layer is processed through a Spatial Pyramid Recalibration (SPR) module (Yu et al., 2024), which is designed to reinforce the important information in the feature map $Fi$. This allows the model to focus more on the critical regions of the feature map while ignoring irrelevant noise, thereby improving the precision of anomaly region segmentation. The processed feature maps $F_{i}^{\prime}$ are then concatenated together to form a new feature map $F^{\prime}\in\mathbb{R}^{H\times W\times 3C}$, as expressed by:
(6) $F^{\prime}=\text{Concat}(F_{1}^{\prime},F_{2}^{\prime},F_{3}^{\prime})$
The concatenated feature map $F^{\prime}$ is passed through a Softmax layer to generate a weight map $W\in\mathbb{R}^{1\times 1\times 3C}$, with the operation for obtaining W as follows:
(7) $W=\text{Softmax}(F^{\prime})$
@@
(8) $T=W\odot\text{Concat}(F_{1},F_{2},F_{3})$
Finally, a channel sum operation is applied to $T$ to convert it back to the original number of channels $C$, producing the output feature map $Y\in\mathbb{R}^{H\times W\times C}$:
(9) $Y=\text{channel\_sum}(T)$
@@
3.3. Segmentation Network
In the works of (Salehi et al., 2020), anomaly scores are typically represented by directly summing the cosine distances of multi-layer features. However, if the discriminative abilities of the features across layers vary, this direct summation may result in suboptimal performance. To address this issue, we introduce a segmentation network that provides additional supervisory signals to guide the feature fusion process. Specifically, we freeze the weights of both the student and teacher networks to train the segmentation network. The synthetic anomalous images are passed as input to both the teacher-student network, while the corresponding binary anomaly masks are used as target labels during training.
In this process, we calculate the similarity between the feature maps $(T_{1},PA\_SD^{1})$, $(T_{2},PA\_SD^{2})$, $(T_{3},PA\_SD^{3})$, as shown in Equation (2), and the results are upsampled to a quarter of the input image size, i.e., the size of $X_{1}$. After upsampling, the resulting feature maps $\widehat{X}_{1}$, $\widehat{X}_{2}$, $\widehat{X}_{3}$ are concatenated into $\widehat{X}$ and passed to the segmentation network for further processing. The segmentation network mainly consists of two PA_Residual Blocks and a Rectangular Self-Calibration Module (RCM) (Ni et al., 2024). In (Zhang et al., 2022), the segmentation network employs the Atrous Spatial Pyramid Pooling (ASPP) module to increase the receptive field without introducing additional computational costs. However, the dilated convolutions used by ASPP tend to overlook small foreground objects in the image. This limitation is particularly problematic in industrial anomaly detection tasks, such as those involving datasets like MVTec AD (Bergmann et al., 2019), where most anomalies are small region anomalies. Consequently, the use of ASPP can lead to over-segmentation, where the detected anomaly regions include many non-anomalous parts, thereby reducing the model’s ability to accurately localize anomalies. To address this issue, we replace the ASPP module with the RCM proposed in (Ni et al., 2024). This adjustment allows the segmentation network to focus more on the foreground objects by adjusting the rectangular attention regions, improving the localization of foreground objects and precisely segmenting anomalous regions. This ensures that the anomaly regions are accurately delineated while reducing the segmentation of non-anomalous areas.
Table 1. Average Image-Level Anomaly Detection AUC (%) Across All Categories on the MVTec AD Dataset
STPM(Wang et al., 2021) CutPaste(Li et al., 2021) DRAEM(Zavrtanik et al., 2021) DSR(Zavrtanik et al., 2022) PatchCore(Roth et al., 2021) DeSTSeg(Zhang et al., 2022) Ours
95.1 95.2 98 98.2 98.5 98.4 98.9
@@ 3.3. Segmentation Network
(10) $L_{focal}=-\frac{1}{H_{1}W_{1}}\sum_{i,j=1}^{H_{1},W_{1}}(1-q_{ij})^{\gamma}\log(q_{ij})$
@@
Finally, the total segmentation loss is the weighted sum of the focal loss and L1 loss:
(12) $L_{seg}=L_{focal}+L_{l1}$
Table 2. AUC/AP (%) Results for Pixel-Level Anomaly Localization on the MVTec AD Dataset
Category STPM(Wang et al., 2021) CutPaste(Li et al., 2021) DRAEM(Zavrtanik et al., 2021) DSR(Zavrtanik et al., 2022) PatchCore(Roth et al., 2021) DeSTSeg(Zhang et al., 2022) Ours
texture carpet 99.1 / 65.3 98.3 / - 96.2 / 64.4 - / 78.2 99.1 / 66.7 93.5 / 62.9 96.6 / 75.4
grid 99.1 / 45.4 97.5 / - 99.5 / 56.8 - / 68.0 98.9 / 41.0 99.2 / 61.9 98.6 / 58.6
leather 99.2 / 42.9 99.5 / - 98.9 / 69.9 - / 62.5 99.4 / 51.0 99.8 / 77.3 99.7 / 76.4
tile 96.6 / 61.7 90.5 / - 99.5 / 96.9 - / 93.9 96.6 / 59.3 98.0 / 94.4 98.3 / 91.6
wood 95.2 / 47.0 95.5 / - 97.0 / 80.5 - / 68.4 95.1 / 52.3 96.2 / 71.7 97.8 / 82.4
average 97.8 / 52.5 96.3 / - 98.2 / 73.7 - / 74.2 97.8 / 54.1 97.3 / 73.6 98.2 / 76.9
object bottle 98.8 / 80.6 97.6 / - 99.3 / 89.8 - / 91.5 98.9 / 80.1 99.1 / 90.2 99.3 / 91.3
cable 94.8 / 58.0 90.0 / - 95.4 / 62.6 - / 70.4 98.8 / 70.0 97.0 / 59.8 96.6 / 63.8
capsule 98.2 / 35.9 97.4 / - 94.1 / 43.5 - / 53.3 99.1 / 48.1 97.6 / 53.3 99.0 / 59.3
hazelnut 98.9 / 60.3 97.3 / - 99.5 / 88.1 - / 87.3 99.0 / 61.5 99.5 / 89.6 99.6 / 89.3
metal nut 97.2 / 79.3 93.1 / - 98.7 / 91.7 - / 67.5 98.8 / 88.8 98.9 / 93.4 98.9 / 94.2
pill 94.7 / 63.3 95.7 / - 97.6 / 46.1 - / 65.7 98.2 / 78.7 98.4 / 84.9 98.8 / 84.2
screw 98.6 / 26.9 96.7 / - 99.7 / 71.5 - / 52.5 99.5 / 41.4 98.5 / 59.1 97.6 / 58.4
toothbrush 98.9 / 48.8 98.1 / - 98.1 / 54.7 - / 74.2 98.9 / 51.6 99.4 / 76.9 99.4 / 71.5
transistor 81.9 / 44.4 93.0 / - 90.0 / 51.7 - / 41.1 96.2 / 63.2 86.1 / 79.1 94.2 / 65.8
zipper 98.0 / 54.9 99.3 / - 98.6 / 72.3 - / 78.5 99.0 / 64.0 99.1 / 83.9 99.2 / 84.0
average 96.0 / 55.2 95.9 / - 97.2 / 67.1 - / 68.2 98.6 / 64.8 97.4 / 77.0 98.3 / 76.2
average 96.6 / 54.3 96.0 / - 97.5 / 69.3 - / 70.2 98.4 / 61.2 97.4 / 75.9 98.2 / 76.4
4. Experiments
4.1. Dataset
We conducted a comprehensive evaluation of the proposed method using the MVTec AD (Bergmann et al., 2019) dataset. It is one of the most widely used benchmarks in the field of anomaly detection and localization, consisting of a total of 15 categories, including 10 object categories and 5 texture categories. The training set for each category contains several hundred normal images, while the test set consists of a mix of normal and anomalous images. Additionally, each anomalous image in the test set is provided with ground truth annotations for performance evaluation. The image resolutions range from $700\times 700$ to $1024\times 1024$ pixels. For the anomalous source image $A$ as shown in Equation (1), we used images from the Describable Textures Dataset (DTD) (Saito and Rehmsmeier, 2015). Although (Zavrtanik et al., 2021) shows that other datasets, such as ImageNet, can achieve comparable performance, the DTD dataset is more compact, easier to use, and better suited for generating anomalies.
4.2. Evaluation Metrics
We used three levels of evaluation metrics for the examination of model’s detection performance. For image-level evaluation, we adopt the area under the ROC curve (AUC), which is a common practice in the anomaly detection field. Rather than AUC, Average Precision (AP), as highlighted in (Saito and Rehmsmeier, 2015), is reported for the pixel-level evaluation considering the significant class imbalance in the dataset. In practical applications, such as industrial defect detection and medical imaging lesion detection, users are more concerned with whether the model can locate the anomalous instances entirely or partially, instead of detecting every pixel precisely. In such case, the Region Overlap Ratio (PRO) metric proposed in (Bergmann et al., 2020) is used for the instance-level evaluation. It assigns equal weights to connected regions of different sizes in the ground truth annotations and calculates the overlap between the predicted and ground truth regions within a user-defined false positive rate (e.g., $30\%$). However, since instance recall is particularly important in real-world scenarios, we propose using Instance Average Precision (IAP) as a more intuitive evaluation metric. We define anomalous instances as the largest connected regions in the ground truth. In the predicted results, an anomalous instance is considered successfully detected if more than $50\%$ of the pixels in the instance are correctly predicted as positive. By setting different thresholds, we can plot the curve of pixel-level precision versus instance-level recall, and the area under this curve represents the IAP. Furthermore, for applications requiring high recall, we also calculate the precision at a recall rate of $k\%$ and denote it as IAP@$k$. In the experiments, to simulate high-risk scenarios, we set $k$ to 90.
Figure 5. Visual comparison of anomaly segmentation between our method and DeSTSeg. For each example, first column: input image; second column: Ground Truth; third and fourth columns: predicted segmentation results.
@@ 4.2. Evaluation Metrics
Table 3. Instance-Level Anomaly Detection IAP/IAP@90 (%) Results on the MVTec AD Dataset
Category STPM(Wang et al., 2021) DRAEM(Zavrtanik et al., 2021) PatchCore(Roth et al., 2021) DeSTSeg(Zhang et al., 2022) Ours
@@
4.3. Implementation Details
In the experiments, the teacher network is based on a ResNet18 (He et al., 2016) model pre-trained on ImageNet (Deng et al., 2009), with the weights of the teacher network remaining fixed throughout the training process. The encoder of the denoised student network adopts an architecture that improved the residual blocks of ResNet18, while the decoder also adopts the reverse architecture of the improved ResNet18 residual blocks (replacing all downsampling with bilinear upsampling). Both the encoder and decoder are constructed by concatenating four residual blocks in an equal manner, with all weights randomly initialized and involved in training. The segmentation network consists of two improved residual blocks(PA_Residual Blocks), and a Rectangular Self-Calibration module (RCM), with all weights also randomly initialized and involved in the training process. Both the denoised student network and the segmentation network are trained using Stochastic Gradient Descent (SGD). The learning rate for the denoised student network is set to 0.5, the learning rate for the improved residual blocks in the segmentation network is set to 0.1, and the learning rate for the RCM module is set to 0.01, with a batch size of 16. During training, the denoised student network is first trained for 3000 iterations, after which its weights are fixed. Then, the segmentation network undergoes an additional 4000 iterations of training. All input images are resized to $256\times 256$ during both training and testing.
4.4. Results
We primarily compared our method with current state-of-the-art approaches using the evaluation metrics mentioned in Sec. 4.2. Experimental results that were unavailable are marked as ”-”. Below, we will primarily focus on comparing our method with others in terms of image-level anomaly detection, pixel-level anomaly localization, and instance-level anomaly detection, while also presenting visual comparison results that are closest to our method.
@@ 4.4. Results
For the image-Level anomaly detection, we present the AUC scores of our method and other methods in the image-level anomaly detection task in Tab. 1, with a focus on the average results across the 15 categories of the MVTec AD dataset. As observed, our method outperforms others, achieving an Image_AUC of $98.9\%$. In Tab. 2, we report the AUC and AP metrics for the pixel-level anomaly localization task. Our method achieves nearly identical AUC scores compared to PatchCore (Roth et al., 2021), while setting a new benchmark in terms of AP. Additionally, our approach demonstrates substantial improvements across most categories, with scores either at or near the highest, strongly suggesting that it performs robustly across a wide range of categories.
@@
For the instance-level anomaly detection, the experimental results are shown in Tab. 3, where we report the IAP and IAP@90 metrics to comprehensively assess the model’s performance. The results indicate that our method achieves state-of-the-art performance on both metrics, demonstrating its exceptional anomaly detection capabilities. Notably, our method reaches an average IAP@90 of $62.7\%$. This means that when the model successfully detects $90\%$ of the anomalous instances, the pixel-level precision reaches $62.7\%$. In other words, this corresponds to only a $37.3\%$ pixel-level false positive rate, which is relatively low and indicates a high level of robustness and reliability in many real-world applications. The high precision and low false positive rate of our method, especially in complex and high-risk industrial environments, provide a strong guarantee for practical applications. This makes it particularly valuable in industrial domains and other critical areas.
In addition, we provide a visual comparison between our method and the DeSTSeg approach (Zhang et al., 2022). As shown in Fig. 5, the segmentation results from our method clearly outperform those of DeSTSeg in terms of visual quality. Notably, for the five texture categories on the right side, our approach excels in accurately focusing on the abnormal regions while effectively minimizing irrelevant noise.

### v4

@@ 3.2 Parallel Convolutional Attention Recalibration Module
ResNet18 employs residual blocks within its convolutional layers (i.e., conv2_x, conv3_x, conv4_x, conv5_x). We propose a novel Parallel Convolutional Attention Recalibration (PCAR) Module to enhance these blocks. Employing parallel convolutions to extract diverse spatial information, the PCAR module aims to improve the localization and segmentation of anomalous regions while suppressing irrelevant noise. Specifically, PCAR replaces the second $3\times 3$ convolution layer within the standard ResNet18 residual block. The detailed architecture of the PCAR module is illustrated in Fig. 4.
@@
The PCAR module takes the feature map $F\in\mathbb{R}^{H\times W\times C}$ (output from the block’s first $3\times 3$ convolution) as input. Here, $H,W,C$ denote height, width, and channels, respectively. This input $F$ is processed by three parallel $3\times 3$ convolutional layers. Each path independently extracts features ($F_{i}$, for $i=1,2,3$), thus enhancing the diversity of the features. This operation is defined as:
$F_{i}=\text{Conv}{3\times 3}(F)$ (5)
Each resulting feature map $F_{i}$ is then processed by a Spatial Pyramid Recalibration (SPR) module [22], designed to reinforce salient information within $F_{i}$. The SPR module helps focus on critical feature regions while suppressing noise, thus enhancing segmentation precision. The SPR-processed feature maps ($F_{i}^{\prime}$) are subsequently concatenated along the channel dimension:
$F^{\prime}=\text{Concat}(F_{1}^{\prime},F_{2}^{\prime},F_{3}^{\prime})$ (6)
Channel weights $W\in\mathbb{R}^{1\times 1\times 3C}$ are generated from $F^{\prime}$ via Softmax activation:
$W=\text{Softmax}(F^{\prime})$ (7)
@@
$T=W\odot\text{Concat}(F_{1},F_{2},F_{3})$ (8)
Finally, a channel reduction operation is applied to $T$ to restore the original channel dimension $C$, producing the final output feature map $Y\in\mathbb{R}^{H\times W\times C}$:
$Y=\text{channel\_sum}(T)$ (9)
@@
Figure 4: Overall architecture of the PCAR module. The feature map is represented by its dimensions $H$, $W$, and $C$, denoting height, width, and the number of channels, respectively. The Softmax operation recalibrates channel attention weights to capture long-range channel dependencies. The symbol $\otimes$ indicates element-wise multiplication, and channel_sum refers to summation along the channel dimension.
3.3 Segmentation Network
In the works of [15], anomaly scores are typically represented by directly summing the cosine distances of multi-layer features. However, if the discriminative abilities of the features across layers vary, this direct summation may result in suboptimal performance. To address this issue, we introduce a segmentation network that provides additional supervisory signals to guide the feature fusion process. Specifically, we freeze the weights of both the student and teacher networks to train the segmentation network. The synthetic anomalous images are passed as input to both the teacher-student network, while the corresponding binary anomaly masks are used as ground truth during training.
For input to the segmentation network, we compute the cosine similarity maps $X_{1},X_{2},X_{3}$ between the paired features $(T_{1},PA\_SD^{1})$, $(T_{2},PA\_SD^{2})$, $(T_{3},PA\_SD^{3})$ using Equation (2). These similarity maps are then upsampled to match the spatial dimensions of $X_{1}$. The upsampled maps, denoted $\widehat{X}_{1},\widehat{X}_{2},\widehat{X}_{3}$, are concatenated along the channel dimension to form the input $\widehat{X}$ for the segmentation network. Our segmentation network primarily consists of two enhanced PA_Residual Blocks and a Rectangular Self-Calibration Module (RCM) [11]. While prior work like [25] used Atrous Spatial Pyramid Pooling (ASPP) for multi-scale context, its dilated convolutions can overlook small objects, leading to imprecise localization and over-segmentation, especially for the small anomalies common in datasets like MVTec AD [2]. Therefore, we replace ASPP with RCM [11]. RCM utilizes adjustable rectangular attention regions, enabling better focus on foreground anomalies of varying shapes and improving localization precision. This helps accurately delineate anomaly boundaries while minimizing the inclusion of non-anomalous areas.
@@ 3.3 Segmentation Network
$L_{focal}=-\frac{1}{H_{1}W_{1}}\sum_{i,j=1}^{H_{1},W_{1}}(1-q_{ij})^{\gamma}\log(q_{ij})$ (10)
@@
Finally, the total segmentation loss is the sum of the focal loss and L1 loss:
$L_{seg}=L_{focal}+L_{l1}$ (12)
@@
Table 1: Average Image-Level Anomaly Detection AUC (%) Across All Categories on the MVTec AD Dataset
STPM[18] DRAEM[23] DSR[24] PatchCore[13] DeSTSeg[25] CND[19] Ours
95.1 98 98.2 98.5 98.4 98.6 98.9
4 Experiments
4.1 Dataset
We comprehensively evaluated our proposed method on the MVTec AD [2] dataset. MVTec AD is a widely recognized benchmark for unsupervised anomaly detection and localization, consisting of a total of 15 categories, including 10 object categories and 5 texture categories. Each category provides a training set of several hundred normal images and a test set containing both normal and anomalous images. Anomalous images in the test set include pixel-precise ground truth annotations for evaluation. Image resolutions range from $700\times 700$ to $1024\times 1024$ pixels. For the external data source $A$ used in synthetic anomaly generation (Equation (1)), we utilized the Describable Textures Dataset (DTD) [5]. Although [23] shows that other datasets, such as ImageNet, can achieve comparable performance, the DTD dataset is more compact, easier to use, and better suited for generating anomalies.
4.2 Evaluation Metrics
We used three levels of evaluation metrics for the examination of model’s detection performance. For image-level evaluation, we use the Area Under the ROC Curve (AUC), a standard metric in anomaly detection. At the pixel level, we use Average Precision (AP) [14], as it better handles the significant class imbalance typical in anomaly detection datasets compared to AUC. For instance-level evaluation, practical applications (e.g., industrial defect and medical lesion detection) often prioritize locating anomalous instances (entirely or partially) over precise pixel-wise detection. One common metric is the Per-Region Overlap (PRO) score proposed in [3]. PRO assigns equal weight to ground truth connected regions of varying sizes and measures overlap with predictions at a user-defined false positive rate (e.g., $30\%$). However, recognizing the critical role of instance recall in real-world applications, we propose Instance Average Precision (IAP) as a more intuitive metric. For IAP, anomalous instances are defined as the largest connected regions in the ground truth. An instance is considered detected if over $50\%$ of its pixels are correctly classified as positive. By setting different thresholds, we can plot the curve of pixel-level precision versus instance-level recall, and the area under this curve represents the IAP. Furthermore, for scenarios demanding high recall, we also compute precision at an instance recall of $k\%$ (IAP@$k$). In our experiments, $k=90$ to simulate high-risk applications.
Table 2: AUC/AP (%) Results for Pixel-Level Anomaly Localization on the MVTec AD Dataset
Category STPM[18] DRAEM[23] DSR[24] PatchCore[13] DeSTSeg[25] CND[19] Ours
texture carpet 99.1 / 65.3 96.2 / 64.4 - / 78.2 99.1 / 66.7 93.5 / 62.9 99.3 / 70.7 96.6 / 75.4
grid 99.1 / 45.4 99.5 / 56.8 - / 68.0 98.9 / 41.0 99.2 / 61.9 98.4 / 25.8 98.6 / 58.6
leather 99.2 / 42.9 98.9 / 69.9 - / 62.5 99.4 / 51.0 99.8 / 77.3 99.5 / 50.1 99.7 / 76.4
tile 96.6 / 61.7 99.5 / 96.9 - / 93.9 96.6 / 59.3 98.0 / 94.4 97.7 / 73.4 98.3 / 91.6
wood 95.2 / 47.0 97.0 / 80.5 - / 68.4 95.1 / 52.3 96.2 / 71.7 96.4 / 63.3 97.8 / 82.4
average 97.8 / 52.5 98.2 / 73.7 - / 74.2 97.8 / 54.1 97.3 / 73.6 98.2 / 56.6 98.2 / 76.9
object bottle 98.8 / 80.6 99.3 / 89.8 - / 91.5 98.9 / 80.1 99.1 / 90.2 99.0 / 81.8 99.3 / 91.3
cable 94.8 / 58.0 95.4 / 62.6 - / 70.4 98.8 / 70.0 97.0 / 59.8 98.2 / 64.1 96.6 / 63.8
capsule 98.2 / 35.9 94.1 / 43.5 - / 53.3 99.1 / 48.1 97.6 / 53.3 98.2 / 36.9 99.0 / 59.3
hazelnut 98.9 / 60.3 99.5 / 88.1 - / 87.3 99.0 / 61.5 99.5 / 89.6 98.8 / 53.3 99.6 / 89.3
metal nut 97.2 / 79.3 98.7 / 91.7 - / 67.5 98.8 / 88.8 98.9 / 93.4 95.5 / 68.4 98.9 / 94.2
pill 94.7 / 63.3 97.6 / 46.1 - / 65.7 98.2 / 78.7 98.4 / 84.9 98.8 / 80.2 98.8 / 84.2
screw 98.6 / 26.9 99.7 / 71.5 - / 52.5 99.5 / 41.4 98.5 / 59.1 99.0 / 26.2 97.6 / 58.4
toothbrush 98.9 / 48.8 98.1 / 54.7 - / 74.2 98.9 / 51.6 99.4 / 76.9 99.0 / 49.7 99.4 / 71.5
transistor 81.9 / 44.4 90.0 / 51.7 - / 41.1 96.2 / 63.2 86.1 / 79.1 94.5 / 57.3 94.2 / 65.8
zipper 98.0 / 54.9 98.6 / 72.3 - / 78.5 99.0 / 64.0 99.1 / 83.9 97.6 / 45.1 99.2 / 84.0
average 96.0 / 55.2 97.2 / 67.1 - / 68.2 98.6 / 64.8 97.4 / 77.0 97.8 / 56.3 98.3 / 76.2
average 96.6 / 54.3 97.5 / 69.3 - / 70.2 98.4 / 61.2 97.4 / 75.9 98.0 / 56.4 98.2 / 76.4
4.3 Implementation Details
In the experiment, the teacher network is based on the ResNet18 [9] model pre-trained on ImageNet [8], and the weights of the teacher network remain fixed throughout the training process. The encoder of the denoising student network uses the architecture of PA_ResNet18, and the decoder also uses the reverse architecture of PA_ResNet18 (replacing all downsampling with bilinear upsampling). Both the encoder and decoder are constructed by concatenating four residual blocks in an equal manner, with all weights randomly initialized and participating in training. The segmentation network consists of two improved residual blocks (PA_Residual Blocks) and a rectangular self-calibration module (RCM), all of which are also randomly initialized and participate in the training process. The denoising student network and the segmentation network were trained using stochastic gradient descent (SGD). The learning rate of the denoised student network is set to 0.5, the learning rate of the improved residual block in the segmentation network is set to 0.1, the learning rate of the RCM module is set to 0.01, and the batch size is 16. During training, the denoised student network is first trained for 3000 iterations, after which its weights are fixed. The segmentation network then undergoes an additional 4000 iterations of training. All input images were resized to $256\times 256$ during training and testing.
4.4 Results
We compare our method with several leading contemporary approaches using the evaluation metrics detailed in Sec. 4.2. Unavailable experimental results from other methods are marked as "-".
Table 3: Instance-Level Anomaly Detection IAP/IAP@90 (%) Results on the MVTec AD Dataset
Category STPM[18] DRAEM[23] PatchCore[13] DeSTSeg[25] Ours
@@ 4.4 Results
For the image-level anomaly detection, we present the AUC scores of our method and other methods in the image-level anomaly detection task in Table 1, with a focus on the average results across the 15 categories of the MVTec AD dataset. As observed, our method outperforms others, achieving an Image_AUC of $98.9\%$. In Table 2, we report the AUC and AP metrics for the pixel-level anomaly localization task. Our method achieves nearly identical AUC scores compared to PatchCore [13], while setting a new benchmark in terms of AP. Additionally, our approach demonstrates substantial improvements across most categories, with scores either at or near the highest, strongly suggesting that it performs robustly across a wide range of categories.
@@
Instance-level anomaly detection results are presented in Table 3, featuring the IAP and IAP@90 metrics for comprehensive assessment. The results show that our method achieves highly competitive performance on both metrics, demonstrating strong anomaly detection capabilities. Notably, our method reaches an average IAP@90 of $62.7\%$. This means that when the model successfully detects $90\%$ of anomalous instances, the pixel-level precision is $62.7\%$. This translates to a pixel-level false positive rate of only $37.3\%$ under these conditions, suggesting a good balance of recall and precision, indicative of robustness for real-world applications. The achieved precision at high instance recall suggests our method’s potential utility in demanding industrial environments and other critical areas where reliable anomaly detection is paramount.
Additionally, Fig. 5 provides a visual comparison of our method against the DeSTSeg approach [25]. The qualitative results indicate that our method produces segmentation masks with improved visual quality compared to DeSTSeg. Particularly for the five texture categories (shown on the right), our method demonstrates a better capability to accurately highlight anomalous regions while effectively suppressing irrelevant noise.
Figure 5: Visual comparison of anomaly segmentation between our method and DeSTSeg. For each example, first column: input image; second column: Ground Truth; third and fourth columns: predicted segmentation results.
4.5 Ablation Studies
In Table 4, we conducted ablation experiments to effectively evaluate the proposed three modules. Specifically, we adopted the network architecture from [25] as the baseline and performed a detailed analysis and comparative experiments to assess the independent effects of each module as well as their combined impact on network performance. The experimental results demonstrate that, by comparing Experiment 2 with 5 and Experiment 3 with 6, integrating the proposed PCAR module with either the RCM or AFF module significantly enhances the network’s performance. Furthermore, the results of Experiment 7 further reveal that incorporating the PCAR, RCM, and AFF modules into the network architecture achieves optimal performance. This combination not only enables more accurate detection and localization but also substantially improves the overall network effectiveness, thereby validating the efficiency and superiority of our proposed approach.